OpenStudy (anonymous):

Solve for x: 3(x + 1) = -2(x - 1) - 4. 1 -1 -5 -25

OpenStudy (anonymous):

@♪Emoticon

OpenStudy (anonymous):

Well we use the distributive property first. Do you know how that goes?

OpenStudy (solomonzelman):

expand 3(x + 1) and -2(x - 1) add like terms on the right side subtract 3 from both sides add 2x to both sides divide both sides by 5

OpenStudy (anonymous):

Would the answer be -1?

OpenStudy (anonymous):

I'm not too good with distributive prop.

OpenStudy (solomonzelman):

No, not -1, do what I told you.

OpenStudy (anonymous):

if 5x divided by 5 would it be 1?

OpenStudy (anonymous):

1x, but 1x is just x

OpenStudy (anonymous):

So since x is just 1 it is A correct?

OpenStudy (anonymous):

no, 1x is x. 1x is just x once so its just x alone.

OpenStudy (anonymous):

but wouldn't x alone equal 1 therefore making the answer A?

OpenStudy (anonymous):

no, x alone is x we don't know the value yet

OpenStudy (anonymous):

How do we find out the volume?

OpenStudy (anonymous):

3(x + 1) = -2(x - 1) - 4 3x + 1 = -2x + 1 - 4 Now subtract 4 from 1: 3x + 1 = -2x -3 Subtract 1 from both sides: 3x = -2x - 4

OpenStudy (anonymous):

One last step, do you know what it is?

OpenStudy (anonymous):

distribute?